Edmond blocked 27 of 36 shots based on information what's the probality of him blocking the next 16 shots
Answer by stanbon(75887)

Edmond blocked 27 of 36 shots based on information what's the probality of him blocking the next 16 shots
--------
P(blocked) = 27/36 = 3/4
-----
P(block next 16) = (3/4)^16 = 0.01
----------------------------------------
Cheers,
Stan H.
======================